Identify the solution(s) of the system of equations, if any.
5x+2y=2 and 4x-2y=6

Answer by Alan3354(69443)

Identify the solution(s) of the system of equations, if any.
5x+2y=2 and 4x-2y=6
------------
5x+2y=2
4x-2y=6
---------- Add them